The length of a rectangle box is 3 inches more than twice the width. The perimeter is 48 inches. Find the length and width [Searched for: help me with a word problem
Answer by lwsshak3(11628)

The length of a rectangle box is 3 inches more than twice the width. The perimeter is 48 inches. Find the length and width
**
let x=width of rectangle box
2x+3=length of rectangle box
2*width+2*length=perimeter
..
2x+2(2x+3)=48
x+2x+3=24
3x=21
x=7
2x+3=17
width of rectangle box=7 inches
length of rectangle box=17 inches